# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@ 5.1. Blowing up analysis
(5.10) $\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\leq C\tau^{-2}(\varepsilon_{i}+o(i))+C\tau.$
This will lead to a contradiction of (5.6)${}_{2}$ if we chose sufficient small $\tau_{0}\in(0,\frac{1}{4})$ and sufficient large $i_{0}$ such that for $i\geq i_{0}$, it holds that
(5.11) $\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}|\widehat{\pi}_{i}|^{\frac{3}{2}}\right)^{2}\leq\frac{1}{4}.$
To prove (5.10), let $\eta\in C_{0}^{\infty}(B_{1}(0))$ be a cut-off function of $B_{\frac{3}{8}}(0)$. For any $t$, $-(\frac{3}{8})^{2}\leq t\leq 0$, we define
$\widehat{\pi}_{(1)}^{i}(x,t)=\int_{\mathbb{T}^{3}}\nabla_{x}^{2}G(x-y)\eta(y)\varepsilon_{i}\left[\widehat{\textbf{u}}^{i}\otimes\widehat{\textbf{u}}^{i}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right](y,t){\,\mathrm{d}y},$
@@
(5.12) $-\Delta\widehat{\pi}_{(2)}^{i}=0\text{ in }B_{\frac{3}{8}}(0).$
@@
(5.13) $\left\|\widehat{\pi}_{(1)}^{i}\right\|_{L^{\frac{3}{2}}(\mathbb{T}^{3})}\leq C\varepsilon_{i}\left[\left\|\widehat{\textbf{u}}^{i}\right\|_{L^{3}(B_{1}(0))}^{2}+\left\|\nabla\widehat{{\bf{d}}}^{i}\right\|_{L^{3}(B_{1}(0))}^{2}+\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}\|F({\bf{d}}^{i})\|_{L^{\frac{3}{2}}(B_{1}(0))}\right]\leq C(\varepsilon_{i}+o(i)).$
@@
(5.14) $\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}(|\widehat{\pi}^{i}|^{\frac{3}{2}}+|\widehat{\pi}_{(1)}^{i}|^{\frac{3}{2}})\leq C\tau(1+\varepsilon_{i}+o(i)).$
Combining (5.13) and (5.14) yields (5.10), and the contradiction (5.11) is achieved. ∎
5.2. An almost boundedness result
By iterating Lemma 5.2 and utilizing the Riesz potential in Morrey spaces (see [1, 27]) we can show the following local $L^{p}$ estimate for all $p>1$:
Lemma 5.3.
For any $M>0$, there exists $\varepsilon_{0}>0$ depending on $M$, such that for a suitable weak solution $({{\bf{v}}},{\pi},{{\bf{d}}})$ of (3.3) in ${\mathbb{T}^{3}}\times(0,\infty)$, if for $z_{0}=(x_{0},t_{0})\in{\mathbb{T}^{3}}\times(r_{0}^{2},\infty)$, and $r_{0}>0$, it holds that $|{\bf{z}}|\leq M,|{{\bf{d}}}|\leq M$ in $\mathcal{Q}_{r_{0}}(z_{0})$, and $\Theta_{z_{0}}(r_{0})\leq\varepsilon_{0}^{3}$, then for any $1<p<\infty$, it holds true that
@@ 5.2. An almost boundedness result
From the assumption, we have that $\Phi_{z}(r_{0}/2)\leq 8\varepsilon_{0}^{3}$ holds for any $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$. Then we apply Lemma 5.2 repeatedly on $\mathcal{Q}_{\frac{r_{0}}{2}}(z)$, and we have that for any $k\geq 1$ and $\tau_{0}\in(0,1/4)$,
@@
(5.15) $s^{-2}\int_{\mathcal{Q}_{s}(z)}(|{\bf{v}}|^{2}+|\nabla{{\bf{d}}}|^{3}+|\pi|^{\frac{3}{2}})\leq C(1+\varepsilon_{0}^{3})\left(\frac{s}{r_{0}}\right)^{3\theta_{0}},$
@@
(5.16) $s^{-1}\int_{\mathcal{Q}_{s}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
@@
Then (5.15) and (5.16) read
[…]
Now we view (1.1)${}_{3}$ as
@@
where $\mathbb{F}$ satisfies
@@
[…]
@@
[…]
Apply the Riesz potential estimates, we conclude that
(5.17) $\|\nabla{\widecheck{\bf d}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{F}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$
@@
(5.18) $\|\nabla({{\bf{d}}}-{\widecheck{\bf d}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{2}(z_{0})})}\leq C(p,r_{0},\varepsilon_{0}).$
@@
(5.19) $\left\{\begin{array}[]{l}\partial_{t}\widecheck{{\bf{v}}}-\Delta\widecheck{{\bf{v}}}+\nabla\widecheck{\pi}\\ =-\nabla\cdot[\eta^{2}({{\bf{u}}}\otimes{{\bf{u}}})+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}}){\rm I}_{3}]=:-\mathop{\rm div}\nolimits\mathsf{X},\\ \nabla\cdot\widecheck{{\bf{v}}}=0,\\ \widecheck{{\bf{v}}}(\cdot,0)=0.\end{array}\right.$
By using the Oseen kernel, we can estimate $\widecheck{{\bf{v}}}$ by
$|\widecheck{{\bf{v}}}(x,t)|\leq C\mathcal{I}_{1}(|\mathsf{X}|)(x,t).$
@@
(5.20) $\|\widecheck{{\bf{v}}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{X}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$
@@
(5.21) $\|{{\bf{v}}}-\widecheck{{\bf{v}}}\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,r_{0},\varepsilon_{0}).$
Bootstrapping $\alpha$ in the previous estimates (5.17), (5.18), (5.20) and (5.21) to get $\alpha\uparrow\frac{1}{2}$, and then by the embedding theorem for the Morrey spaces, we obtain that
@@
$-\Delta{\pi}=\mathop{\rm div}\nolimits^{2}[{{\bf{u}}}\otimes{{\bf{u}}}+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}})].$
5.3. The A-B-C-D Lemma and Criteria for the regular points
@@
Proposition 5.4.
If $z_{0}$ satisfies the assumption in Lemma 5.2, then $z_{0}$ is a regular point.
@@
By Lemma 5.3, we have that $({{\bf{u}}},{\pi},\nabla{{\bf{d}}})\in L^{p}(\mathcal{Q}_{r_{0}/4}(z_{0}))$ for $1<p<\infty$. Applying the regularity estimate for the generalized Stokes system (c.f. [28, Lemma 2.2]) to the ${\bf{v}}$ equation in (3.3)${}_{1}$, we get that ${\bf{v}}\in C^{\alpha}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0}))$ for some $\alpha\in(0,1)$. On the other hand, the ${{\bf{d}}}$ equation reads
@@
In fact we can improve Proposition 5.4 into the following theorem:
Theorem 5.5.
There exists $\varepsilon_{1}>0$ such that if $({{\bf{u}}},P,{{\bf{d}}})$ is a suitable weak solution to (1.1), which satisfies, for $z_{0}\in\mathbb{T}^{3}\times(0,\infty)$,
(5.22) $\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}(z_{0})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\leq\varepsilon_{1}^{2},$
@@
To prove Theorem 5.5, we will need some preparations. The first step is to note that we have the following interpolation lemma.
Lemma 5.6.
@@
(5.23) $\int_{B_{r}(0)}|v|^{q}(x,t){\,\mathrm{d}x}\lesssim\left(\int_{B_{r}(0)}|\nabla v|^{2}(x,t){\,\mathrm{d}x}\right)^{\frac{q}{2}-a}\left(\int_{B_{r}(0)}|v|^{2}(x,t)\right)^{a}$
@@
With application of Lemma 5.6 we can establish the following result.
Lemma 5.7.
@@
(5.24) $C(r)\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\frac{\rho}{r}\right)^{3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho).$
@@
$\displaystyle\lesssim\int_{B_{r}(0)}\left(\left||{{\bf{v}}}|^{2}-(|{{\bf{v}}}|^{2})\rho\right|+\left||\nabla{{\bf{d}}}|^{2}-(|\nabla{{\bf{d}}}|^{2})_{\rho}\right|\right){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim\rho\int_{B_{\rho}(0)}(|{{\bf{v}}}||\nabla{{\bf{v}}}|+|\nabla{{\bf{d}}}||\nabla^{2}{{\bf{d}}}|){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim\rho^{\frac{3}{2}}\left(\rho^{-1}\int_{B_{\rho}(0)}|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{1}{2}}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}\right){\,\mathrm{d}x}$
$\displaystyle\lesssim\rho^{\frac{3}{2}}A^{\frac{1}{2}}(\rho)\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}+\left(\frac{r}{\rho}\right)^{3}A(\rho).$
Applying Lemma 5.6 with $q=3$, $a=\frac{3}{4}$ we can get
@@
$\displaystyle\lesssim\rho^{\frac{4}{3}}\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}\left(\rho^{-1}\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{2}}$
$\displaystyle\lesssim\rho^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)\left(\int_{B_{r}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{2}}$
$\displaystyle\lesssim\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)+\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho),$
@@
Lemma 5.8.
Under the same assumption of Lemma 5.7, it holds that for any $2\displaystyle 0<r<\rho/$ 2,
(5.25) $D(r)\lesssim\frac{r}{\rho}D(\rho)+\left(\frac{\rho}{r}\right)^{2}\left[A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)+\rho^{\frac{3}{2}}B^{\frac{3}{4}}(\rho)+\rho^{3}\right].$

### v2

@@ 5.1 Blowing up analysis
$\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}^{i}|^{\frac{3}{2}}\leq C\tau^{-2}(\varepsilon_{i}+o(i))+C\tau.$ (58)
This will lead to a contradiction of (50)2 if we chose sufficient small $\tau_{0}\in(0,\frac{1}{4})$ and sufficient large $i_{0}$ such that for $i\geq i_{0}$, it holds that
$\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}\left(|\widehat{{\bf{v}}}^{i}|^{3}+|\nabla\widehat{{\bf{d}}}^{i}|^{3}\right)+\left(\tau_{0}^{-2}\int_{\mathcal{Q}_{\tau_{0}}(0)}|\widehat{\pi}_{i}|^{\frac{3}{2}}\right)^{2}\leq\frac{1}{4}.$ (59)
To prove (58), let $\eta\in C_{0}^{\infty}(B_{1}(0))$ be a cut-off function of $B_{\frac{3}{8}}(0)$. For any $t$, $-(\frac{3}{8})^{2}\leq t\leq 0$, we define
$\displaystyle\widehat{\pi}_{(1)}^{i}(x,t)$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}\nabla_{x}^{2}G(x-y)\eta(y)\varepsilon_{i}\left[{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}\otimes{\left(\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}\right)}+\nabla\widehat{{\bf{d}}}^{i}\odot\nabla\widehat{{\bf{d}}}^{i}-\frac{1}{2}|\nabla\widehat{{\bf{d}}}^{i}|^{2}{\rm I}_{3}-\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}F({{\bf{d}}}^{i}){\rm I}_{3}\right](y,t){\,\mathrm{d}y},$
@@
$\Delta\widehat{\pi}_{(2)}^{i}=0\text{ in }B_{\frac{3}{8}}(0).$ (60)
@@
$\displaystyle\left\|\widehat{\pi}_{(1)}^{i}\right\|_{L^{\frac{3}{2}}(\mathbb{T}^{3})}$ $\displaystyle\leq C\varepsilon_{i}\left[\left\|{\widehat{{\bf{z}}}^{i}+\widehat{{\bf{v}}}^{i}}\right\|_{L^{3}(B_{1}(0))}^{2}+\left\|\nabla\widehat{{\bf{d}}}^{i}\right\|_{L^{3}(B_{1}(0))}^{2}+\frac{r_{i}^{2}}{\varepsilon_{i}^{2}}\|F({\bf{d}}^{i})\|_{L^{\frac{3}{2}}(B_{1}(0))}\right]$
$\displaystyle\leq C(\varepsilon_{i}+o(i)).$ (61)
@@
$\displaystyle\tau^{-2}\int_{\mathcal{Q}_{\tau}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}$ $\displaystyle\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}|\widehat{\pi}_{(2)}^{i}|^{\frac{3}{2}}$
$\displaystyle\leq C\tau\int_{\mathcal{Q}_{\frac{1}{3}}(0)}(|\widehat{\pi}^{i}|^{\frac{3}{2}}+|\widehat{\pi}_{(1)}^{i}|^{\frac{3}{2}})$
$\displaystyle\leq C\tau(1+\varepsilon_{i}+o(i)).$ (62)
Combining (61) and (62) yields (58), and the contradiction (59) is achieved. ∎
5.2 An almost boundedness result
By iterating Lemma 5 and utilizing the Riesz potential in Morrey spaces (see [1, 47]) we can show the following local $L^{p}$ estimate for all $p>1$:
Lemma 6.
For any $M>0$, there exists $\varepsilon_{0}>0$ depending on $M$, such that for a suitable weak solution $({{\bf{v}}},{\pi},{{\bf{d}}})$ of (11) in ${\mathbb{T}^{3}}\times(0,\infty)$, if, for $z_{0}=(x_{0},t_{0})\in{\mathbb{T}^{3}}\times(r_{0}^{2},\infty)$, and $r_{0}>0$, it holds that $|{\bf{z}}|\leq M,|{{\bf{d}}}|\leq M$ in $\mathcal{Q}_{r_{0}}(z_{0})$, and $\Theta_{z_{0}}(r_{0})\leq\varepsilon_{0}^{3}$, then for any $1<p<\infty$, it holds true that
@@ 5.2 An almost boundedness result
From the assumption, we have that $\Phi_{z}(r_{0}/2)\leq 8\varepsilon_{0}^{3}$ holds for any $z\in\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})$. Then we apply Lemma 5 repeatedly on $\mathcal{Q}_{\frac{r_{0}}{2}}(z)$, and we have that for any $k\geq 1$ and $\tau_{0}\in(0,1/4)$,
@@
$s^{-2}\int_{\mathcal{Q}_{s}(z)}(|{\bf{v}}|^{2}+|\nabla{{\bf{d}}}|^{3}+|\pi|^{\frac{3}{2}})\leq C(1+\varepsilon_{0}^{3})\left(\frac{s}{r_{0}}\right)^{3\theta_{0}},$ (63)
@@
$s^{-1}\int_{\mathcal{Q}_{s}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$ (64)
@@
Then (63) and (64) read
$\left\{\begin{aligned} &({{\bf{v}}},\nabla{{\bf{d}}})\in M^{3,3(1-\alpha)}\left(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})\right),\\ &\pi\in M^{\frac{3}{2},3(1-\alpha)}\left(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})\right),\\ &(\nabla{{\bf{v}}},\nabla^{2}{{\bf{d}}})\in M^{2,4-2\alpha}(\mathcal{Q}_{\frac{r_{0}}{2}}(z_{0})).\end{aligned}\right$.
Now we view (4)3 as
@@
where $\mathsf{F}$ satisfies
@@
$|\nabla\widecheck{\bf d}(x,t)|\leq\int_{0}^{t}\int_{\mathbb{R}^{3}}|\nabla\Gamma(x-y,t-s)||\mathsf{F}(y,s)|dyds\leq C\mathcal{I}_{1}(|\mathsf{F}|)(x,t),$ (65)
@@
$I_{\beta}(g)(x,t):=\int_{\mathbb{R}^{4}}\frac{|g(y,s)|}{\delta_{p}^{5-\beta}((x,t),(y,s))}dyds.$
Applying the Riesz potential estimates (for a reference on the Riesz potential estimates, see e.g. [45]), we conclude that
$\|\nabla{\widecheck{\bf d}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{F}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$ (66)
@@
$\|\nabla({{\bf{d}}}-{\widecheck{\bf d}})\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{2}(z_{0})})}\leq C(p,r_{0},\varepsilon_{0}).$ (67)
@@
$\left\{\begin{array}[]{l}\partial_{t}\widecheck{{\bf{v}}}-\Delta\widecheck{{\bf{v}}}+\nabla\widecheck{\pi}\\ =-\nabla\cdot[\eta^{2}({{\bf{u}}}\otimes{{\bf{u}}})+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}}){\rm I}_{3}]\\ =:-\mathop{\rm div}\nolimits\mathsf{X},\\ \\ \nabla\cdot\widecheck{{\bf{v}}}=0,\\ \widecheck{{\bf{v}}}(\cdot,0)=0.\end{array}\right$. (68)
Now we apply the Oseen kernel for estimations of $\widecheck{{\bf{v}}}$; for references about the Oseen kernel, we refer the readers to e.g. [78, 59, 58] Particularly, by using the decaying property of the Oseen kernel (see Proposition 11.1 in [58]), then with a computation similar to (65), we can arrive at
$|\widecheck{{\bf{v}}}(x,t)|\leq C\int_{0}^{t}\int_{\mathbb{R}^{3}}\frac{|{\mathsf{X}}(y,s)|}{\delta_{p}^{4}((x,t),(y,s))}dyds\leq C\mathcal{I}_{1}(|\mathsf{X}|)(x,t).$ (69)
@@
$\|\widecheck{{\bf{v}}}\|_{M^{\frac{3(1-\alpha)}{1-2\alpha},3(1-\alpha)}(\mathbb{R}^{4})}\leq C\|\mathsf{X}\|_{M^{\frac{3}{2},3(1-\alpha)}(\mathbb{R}^{4})}\leq C(1+\varepsilon_{0}).$ (70)
@@
$\|{{\bf{v}}}-\widecheck{{\bf{v}}}\|_{L^{p}(\mathcal{Q}_{\frac{r_{0}}{4}}(z_{0}))}\leq C(p,r_{0},\varepsilon_{0}).$ (71)
Bootstrapping $\alpha$ in the previous estimates (66), (67), (70) and (71) to get $\alpha\uparrow\frac{1}{2}$, and then by the embedding theorem for the Morrey spaces, we obtain that
@@
$\Delta{\pi}=\mathop{\rm div}\nolimits^{2}[{{\bf{u}}}\otimes{{\bf{u}}}+(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3})-(F({{\bf{d}}})-(F({{\bf{d}}}))_{z_{0},r_{0}})].$
5.3 The A-B-C-D Lemma and Criteria for the regular points
@@
Proposition 7.
If $z_{0}$ satisfies the assumption in Lemma 5, then $z_{0}$ is a regular point.
@@
By Lemma 6, we have that $({{\bf{u}}},{\pi},\nabla{{\bf{d}}})\in L^{p}(\mathcal{Q}_{r_{0}/4}(z_{0}))$ for $1<p<\infty$. Applying the regularity estimate for the generalized Stokes system (cf. [51, Lemma 2.2]) to the ${\bf{v}}$ equation in (11)1, we get that ${\bf{v}}\in C^{\alpha}(\mathcal{Q}_{\frac{r_{0}}{8}}(z_{0}))$ for some $\alpha\in(0,1)$. On the other hand, the ${{\bf{d}}}$ equation reads
@@
In fact we can improve Proposition 7 into the following theorem:
Theorem 8.
There exists $\varepsilon_{1}>0$ such that if $({{\bf{u}}},P,{{\bf{d}}})$ is a suitable weak solution to (4), which satisfies, for $z_{0}\in\mathbb{T}^{3}\times(0,\infty)$,
$\limsup_{r\to 0}\frac{1}{r}\int_{\mathcal{Q}_{r}(z_{0})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\leq\varepsilon_{1}^{2},$ (72)
@@
To prove Theorem 8, we will need some preparations. The first step is to note that we have the following interpolation lemma.
Lemma 9.
@@
$\int_{B_{r}(0)}|v|^{q}(x,t){\,\mathrm{d}x}\lesssim\left(\int_{B_{r}(0)}|\nabla v|^{2}(x,t){\,\mathrm{d}x}\right)^{\frac{q}{2}-a}\left(\int_{B_{r}(0)}|v|^{2}(x,t)\right)^{a}$ (73)
@@
With application of Lemma 9 we can establish the following result.
Lemma 10.
@@
$C(r)\lesssim\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho)+\left(\frac{\rho}{r}\right)^{3}A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho).$ (74)
@@
$\displaystyle\lesssim$ $\displaystyle\int_{B_{r}(0)}\left(\left||{{\bf{v}}}|^{2}-(|{{\bf{v}}}|^{2})\rho\right|+\left||\nabla{{\bf{d}}}|^{2}-(|\nabla{{\bf{d}}}|^{2})_{\rho}\right|\right){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim$ $\displaystyle\rho\int_{B_{\rho}(0)}(|{{\bf{v}}}||\nabla{{\bf{v}}}|+|\nabla{{\bf{d}}}||\nabla^{2}{{\bf{d}}}|){\,\mathrm{d}x}+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}$
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{3}{2}}\left(\rho^{-1}\int_{B_{\rho}(0)}|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{1}{2}}$
$\displaystyle+\left(\frac{r}{\rho}\right)^{3}\int_{B_{\rho}(0)}\left(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}\right){\,\mathrm{d}x}$
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{3}{2}}A^{\frac{1}{2}}(\rho)\left(\int_{B_{\rho}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}{\,\mathrm{d}x}\right)^{\frac{1}{2}}+\left(\frac{r}{\rho}\right)^{3}A(\rho).$
Applying Lemma 9 with $q=3$, $a=\frac{3}{4}$ we can get
@@
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{4}{3}}\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}\left(\rho^{-1}\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{4}}$
$\displaystyle+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2}){\,\mathrm{d}x}\right)^{\frac{3}{2}}$
$\displaystyle\lesssim$ $\displaystyle\rho^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)\left(\int_{B_{r}(0)}|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}+r^{-\frac{3}{2}}\left(\int_{B_{r}(0)}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{2}}$
$\displaystyle\lesssim$ $\displaystyle\left(\rho^{\frac{3}{4}}+\frac{\rho^{\frac{9}{4}}}{r^{\frac{3}{2}}}\right)\left(\int_{B_{r}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}A^{\frac{3}{4}}(\rho)+\left(\frac{r}{\rho}\right)^{3}A^{\frac{3}{2}}(\rho),$
@@
Lemma 11.
Under the same assumption of Lemma 10, it holds that for any $0<r<\rho/2$,
$D(r)\lesssim\frac{r}{\rho}D(\rho)+\left(\frac{\rho}{r}\right)^{2}\left[A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho)+\rho^{\frac{3}{2}}B^{\frac{3}{4}}(\rho)+\rho^{3}\right].$ (75)